### v3

# ProCLIP: Progressive Vision-Language Alignment via LLM-based Embedder

Abstract
Contrastive Language-Image Pre-training (CLIP) is constrained by its 77-token limit, lack of multilingual support, and coarse-grained semantic understanding. While replacing CLIP’s text encoder with a Large Language Model (LLM) offers a potential solution, direct contrastive alignment often disrupts the established cross-modal representation due to the lack of alignment priors between the two encoders. In this paper, we present ProCLIP, a curriculum learning framework designed to bridge LLMs and CLIP’s visual space, thereby unlocking CLIP’s potential for long-text, multilingual, and fine-grained understanding. The framework consists of two stages: ❶ Representation Inheritance, which distills CLIP’s original text-space knowledge into the LLM to establish an initial VL prior, and ❷ Contrastive Tuning, which refines the image-text alignment using self-distillation regularization to prevent catastrophic forgetting. To ensure semantic consistency, we introduce instance semantic alignment and embedding structure alignment losses throughout both stages. Extensive experiments demonstrate that ProCLIP improves zero-shot classification by 6.8%–13.5% and exhibits superior performance in long-text, multilingual, and fine-grained cross-modal retrieval tasks. The Code is available at https://github.com/VisionXLab/ProCLIP
ProCLIP: Progressive Vision-Language
Alignment via LLM-based Embedder
Xiaoxing Hu1,2*, Kaicheng Yang3*, Ziyang Gong1, Qi Ming4, Zonhao Guo5, Yu Tian 6, Xiang An3, Ziyong Feng3, Xue Yang1🖂 1Shanghai Jiao Tong University, 2Beijing Institute of Technology, 3DeepGlint 4Beijing University of Technology, 5Tsinghua University Equal contribution  🖂Corresponding author
1 Introduction
CLIP (Radford et al., 2021) establishes robust joint vision-language representations via large-scale contrastive learning, serving as a cornerstone for tasks like retrieval (Yang et al., 2023), generation (Wang et al., 2022), and detection (Wu et al., 2023b). Despite its success, CLIP is constrained by its 77-token English-only supervisory signal (Zhang et al., 2024; Tschannen et al., 2025). Furthermore, the lack of fine-grained textual supervision hampers its ability to capture nuanced semantic details (Hu et al., 2025).
Figure 1: Centered Kernel Alignment (CKA) analysis of visual representations. ProCLIP  outperforms the LLM2CLIP in retaining pretrained knowledge, providing a robust foundation for alignment.
To overcome CLIP’s limitations, Long-CLIP (Zhang et al., 2024) interpolates positional embeddings for longer inputs, yet struggles with fine-grained and multilingual understanding. Meanwhile, FLAME (Cao et al., 2025) and LLM2CLIP (Huang et al., 2024) replace CLIP’s text encoder with LLM-based embedders (BehnamGhader et al., 2024; Lee et al., 2024) to leverage their rich linguistic priors. However, these methods typically perform "from-scratch alignment," directly forcing the image encoder to align with the LLM via contrastive learning. By neglecting CLIP’s original pretrained knowledge, this approach risks overfitting and undermines generalization, especially under data-scarce conditions. This motivates a critical question: How can we systematically leverage CLIP’s pretrained knowledge to achieve efficient cross-modal alignment with LLMs while preserving generalization?
We answer this question by proposing ProCLIP, a progressive alignment framework that reconceptualizes pretrained weights as foundational priors to bridge LLMs and CLIP’s visual space. ProCLIP adopts a two-stage curriculum:
❶ Representation Inheritance: We distill knowledge from the original CLIP text encoder into a lightweight adapter for the LLM. This anchors the frozen LLM to CLIP’s textual space, establishing a robust initial prior.
❷ Contrastive Tuning: Building on this foundation, we refine the joint representation space via contrastive tuning. A self-distillation constraint is applied to the image encoder to safeguard intrinsic vision-language knowledge and mitigate catastrophic forgetting.
We evaluated the preservation of pre-trained knowledge within the vision encoder following alignment, using Centered Kernel Alignment (CKA). As illustrated in Fig. 1, ProCLIP demonstrates superior retention of pre-trained features. This suggests that its carefully designed framework effectively facilitates cross-modal alignment without compromising generalization. ProCLIP  offers several advantages:1) Simplicity: An intuitive framework using pretrained weights as a semantic bridge. 2) Superiority: SOTA performance under identical data constraints. 3) Versatility: Extends CLIP to long-text, multilingual, and fine-grained scenarios with limited data. 4) Potential: A robust paradigm for modular upgrades in multi-encoder models.
Our main contributions are summarized as follows:
  • We identify a key limitation in current LLM-augmented CLIP models: "from-scratch alignment" neglects valuable pretrained priors, leading to compromised generalization.
  • We propose ProCLIP, a progressive framework that distills CLIP’s priors into an LLM-based embedder, followed by self-distilled contrastive tuning to refine alignment without catastrophic forgetting.
  • We demonstrate ProCLIP’s efficacy through extensive experiments. ProCLIP achieves 6.8%–13.5% gains in zero-shot classification and shows significant improvements across short/long-text, multilingual retrieval, and fine-grained understanding tasks.
2 Related Work
Vision-Language Contrastive Learning. Vision-language contrastive learning, exemplified by CLIP (Radford et al., 2021), aims to align multimodal representations in a shared semantic space via large-scale pretraining. As a foundational bridge between modalities, CLIP enables diverse open-vocabulary tasks, including image classification (Zhou et al., 2022b, a; Kim et al., 2024), semantic segmentation (Ding et al., 2022; Li et al., 2022; Ghiasi et al., 2022; Xu et al., 2022; Cho et al., 2024; Lan et al., 2024), and object detection (Du et al., 2022; Kaul et al., 2023). Despite its versatility, CLIP is fundamentally constrained by its text encoder’s fixed input length and limited capacity, which hampers its performance in processing long-form, multilingual, and fine-grained semantics. To address these bottlenecks, Long-CLIP (Zhang et al., 2024) extends input lengths through positional embedding interpolation, while LoTLIP (Wu et al., 2024) aggregates diverse textual information using corner tokens to enhance long-text understanding. However, these methods remain tethered to the original CLIP text encoder’s architecture, preventing the integration of broader open-world knowledge and leaving multilingual support unresolved.
Figure 2: The training pipeline of our proposed ProCLIP. It consists of representation inheritance via cross-architecture distillation and contrastive tuning integrated with self-distillation regularization.
LLM-based Representation Learning. Large Language Models (LLMs) have demonstrated exceptional proficiency across diverse natural language processing tasks (Touvron et al., 2023; Achiam et al., 2023; Bai et al., 2023; Liu et al., 2024a). Recent research has pivoted toward repurposing decoder-only architectures for robust representation learning. Notably, LLM2Vec (BehnamGhader et al., 2024) transforms LLMs into versatile encoders via bidirectional attention and contrastive alignment, while Qwen3-Embedding (Zhang et al., 2025b) leverages massive-scale pretraining to achieve state-of-the-art performance on the MTEB benchmark (Muennighoff et al., 2022). Inspired by these advancements, recent works (Huang et al., 2024; Cao et al., 2025; Zhang et al., 2025a) replace CLIP’s original text encoder with LLM-based embedders to handle multilingual, long-form, and complex textual inputs. However, these methods often employ coarse alignment strategies that inadvertently degrade the model’s inherent generalization. Addressing this gap by developing more refined and efficient alignment techniques remains a critical and open challenge.
Knowledge Distillation. Knowledge distillation (KD) (Hinton et al., 2015) facilitates knowledge transfer from a teacher to a student model, or within a single model via self-distillation (Zhang et al., 2019). In the vision-language domain, several KD techniques have been adapted for CLIP. TinyCLIP (Wu et al., 2023a) employs affinity mimicking to replicate cross-modal interactions, while CLIP-KD (Yang et al., 2024a) integrates feature-based and contrastive distillation to maximize teacher-student similarity. Furthermore, CLIP-CID (Yang et al., 2024b) utilizes cluster-instance discrimination to transfer semantic priors. Unlike these methods that primarily focus on model compression or performance enhancement, our work introduces a self-distillation mechanism specifically designed to mitigate catastrophic forgetting and safeguard the model’s intrinsic generalization during the transition to LLM-based architectures.
3 Methodology
In this section, we first introduce the preliminary (Sec. 3.1), including contrastive language-image pre-training and improving CLIP with an LLM-based embedder. Then we present our proposed ProCLIP framework, which comprises two primary training stages: 1) Representation Inheritance via Cross-Architecture Distillation (Sec. 3.2). 2) Contrastive Tuning Integrated with Self-Distillation Regularization (Sec. 3.3).
3.1 Preliminary
Contrastive Language-Image Pre-training. Contrastive Language-Image Pre-training (CLIP) Radford et al. (2021) learns to align images and text from large-scale image–text pairs through contrastive learning, bridging both modalities in a shared embedding space. Given a batch of image-text pairs $\{(I_{i},T_{i})\}_{i=1}^{\mathcal{B}}$, the image encoder $\mathcal{E}_{I}$ and text encoder $\mathcal{E}_{T}$ map them into the joint semantic space as $\{(v_{i},t_{i})\}_{i=1}^{\mathcal{B}}$. To optimize both encoders in a dual-tower architecture, a symmetric contrastive learning objective is imposed on the resulting representations:
$\begin{split}\mathcal{L}_{CLIP}=-\sum_{i=1}^{\mathcal{B}}\Bigg[&\underbrace{\log\frac{\exp(v_{i}\cdot t_{i}^{\top}/\tau)}{\sum_{j=1}^{\mathcal{B}}\exp(v_{i}\cdot t_{j}^{\top}/\tau)}}_{\text{text-to-image}}+\\ &\underbrace{\log\frac{\exp(t_{i}\cdot v_{i}^{\top}/\tau)}{\sum_{j=1}^{\mathcal{B}}\exp(t_{i}\cdot v_{j}^{\top}/\tau)}}_{\text{image-to-text}}\Bigg].\end{split}$ (1)
However, the native CLIP text encoder is limited to sequences of up to 77 tokens. A common solution is to interpolate the position embeddings of the CLIP text encoder and fine-tune the model. Alternatively, one may replace the CLIP text encoder with an LLM-based embedder. The latter approach not only improves long-text understanding but also enhances multilingual understanding and fine-grained semantic comprehension, resulting in a more versatile vision-language dual-encoder. In this work, we investigate a more efficient alignment strategy that leverages an LLM-based embedder to enhancing CLIP’s comprehensive capabilities.
Improving CLIP with LLM-based Embedder. LLM2CLIP Huang et al. (2024) first introduces an LLM-based embedder into CLIP, demonstrating enhanced long-text understanding. Given an LLM-based encoder $\mathcal{G}_{T}$, it encodes texts $\{T_{i}\}_{i=1}^{\mathcal{N}}$ offline into embeddings $\{t^{\prime}_{i}\}_{i=1}^{\mathcal{N}}$. This process is typically performed in an offline manner. During contrastive fine-tuning, a MLP-based adapter is used to map $\{{t^{\prime}}\}_{i=1}^{\mathcal{N}}$ into the CLIP embedding space for dimensional alignment. The mapped text features and the image features from the CLIP image encoder are then optimized via the contrastive loss in Eq. 1, resulting in a newly aligned representation space. However, applying contrastive learning directly to fine-tuning data to optimize the adapter and vision encoder hinders the convergence of the new dual-tower architecture to an optimal parameter space. This arises because the text representations from the LLM-based embedder and adapter lack prior alignment with the vision encoder. Moreover, unconstrained fine-tuning may also cause excessive drift from the original pre-trained representation, while limited fine-tuning data (e.g., 3M samples) cannot compensate for the knowledge acquired during large-scale pre-training (e.g., 400M samples). To overcome these challenges, we propose a progressive alignment pipeline that improves multimodal alignment while preserving pre-trained knowledge.
3.2 Stage I: Representation Inheritance via Cross-Architecture Distillation.
Given a pre-trained image and text encoder of the CLIP model $\{\mathcal{E}_{I},\mathcal{E}_{T}\}$ and a pre-trained LLM-based embedder $\mathcal{G}_{T}$, our goal is to replace the CLIP text encoder $\mathcal{E}_{T}$ with the LLM-based embedder $\mathcal{G}_{T}$ to enhance comprehensive abilities. Consistent with prior works Huang et al. (2024); Cao et al. (2025); Zhang et al. (2025a), we initially extract embeddings from textual captions offline using $\mathcal{G}_{T}$: $t^{\prime}=\{\mathcal{G}_{T}(T_{i})\in\mathbb{R}^{d}\}_{i=1}^{\mathcal{N}}$, where $d$ represents the dimension of the LLM-based embedder.
The embedding space of the LLM-based embedder exhibits no prior alignment with the CLIP image-text representation space. To bridge this gap, we adopt a cross-architecture distillation strategy that transfers knowledge from the CLIP text embedding space to the LLM embedding space. Specifically, given a batch of texts $\{T_{i}\}_{i=1}^{\mathcal{B}}$, we first utilize a single-layer MLP to unify the dimensions of LLM embeddings and CLIP text embeddings. To facilitate fine-grained semantic alignment, we propose an instance semantic alignment loss, denoted as $\mathcal{L}_{\text{ins}}$. Let $t_{i}^{*}=\text{Adapter}(t^{\prime}_{i})$ and $e_{i}=\mathcal{E}(T_{i})$ represent the projected LLM and original CLIP embeddings, respectively. This loss function leverages text-only data to distill knowledge from CLIP’s text encoder into the LLM-based embedder, defined as follows:
$\mathcal{L}_{\text{ins}}=\sum_{i=1}^{\mathcal{B}}\|t_{i}^{*}-e_{i}\|_{2}.$ (2)
Since $\mathcal{L}_{\text{ins}}$ only focuses on instance-level alignment without capturing the global embedding structure, we propose the embedding structure alignment loss $\mathcal{L}_{\text{struct}}$. This loss measures inter-sample distances within a batch in both the CLIP text encoder and LLM-based embedder spaces, and aligns the two globally by minimizing their pairwise distance discrepancy. The structural loss is:
$\mathcal{L}_{\text{struct}}=\sum_{i<j}^{\mathcal{B}}\|d(t_{i}^{*},t_{j}^{*})-d(e_{i},e_{j})\|_{2},$ (3)
[…]
3.3 Stage II: Contrastive Tuning Integrated with Self-Distillation Regularization.
After the above phase, the $\text{Adapter}(\mathcal{G}_{T})$ has already been preliminarily adapted to CLIP’s vision-language embedding space, making subsequent fine-tuning with vision-language contrastive learning significantly easier. We utilize the InfoNCE loss (Radford et al., 2021) to better align the image embedding $v_{i}$ and the projected LLM embedding $t_{i}^{*}$, which can be formulated as:
$\begin{split}\mathcal{L}_{\text{info}}=-\sum_{i=1}^{\mathcal{B}}\Bigg[&\log\frac{\exp(v_{i}\cdot t_{i}^{*\top}/\tau)}{\sum_{j=1}^{\mathcal{B}}\exp(v_{i}\cdot t_{j}^{*\top}/\tau)}+\\ &\log\frac{\exp(t_{i}^{*}\cdot v_{i}^{\top}/\tau)}{\sum_{j=1}^{\mathcal{B}}\exp(t_{i}^{*}\cdot v_{j}^{\top}/\tau)}\Bigg],\end{split}$ (4)
Method #Data Flickr30k COCO ShareGPT4V Urban-1k DOCCI DCI Avg.
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
Model Architecture: CLIP ViT-B/32
CLIP 400M 80.3 59.8 51.5 30.6 77.3 66.0 60.9 46.8 58.1 53.4 43.1 40.3 61.8 49.5
LLM2CLIP 3M 83.5 70.1 55.6 41.1 94.2 93.4 78.2 84.2 76.2 77.1 62.2 64.4 75.0 71.1
[…]
LLM2CLIP 15M 86.2 72.2 58.5 43.2 95.3 94.2 80.6 85.3 79.2 80.7 64.3 67.6 77.4 73.9
[…]
LLM2CLIP 30M 87.8 72.4 61.1 44.3 96.7 95.9 86.6 88.8 82.9 82.9 67.9 69.5 80.5 75.7
[…]
Model Architecture: CLIP ViT-B/16
CLIP 400M 82.7 63.4 53.7 33.3 76.1 68.9 67.5 53.5 66.8 57.0 45.4 43.0 65.4 45.6
LLM2CLIP 3M 88.0 75.3 60.5 44.8 94.4 94.4 80.6 86.0 81.7 82.2 67.2 69.1 78.7 75.3
[…]
LLM2CLIP 15M 88.9 76.6 62.4 46.5 95.0 95.2 84.5 88.4 83.8 85.1 69.3 72.4 80.7 77.3
[…]
LLM2CLIP 30M 90.2 78.1 65.4 48.5 96.8 96.4 89.7 91.3 86.2 86.8 73.1 74.8 83.6 79.3
[…]
Model Architecture: CLIP ViT-L/14
CLIP 400M 86.6 64.6 57.2 36.4 78.0 68.7 68.4 56.0 65.8 63.1 45.4 43.9 66.9 55.5
LLM2CLIP 3M 92.4 80.1 65.5 49.7 95.2 95.6 83.6 89.0 85.1 85.9 70.0 74.4 82.0 79.1
[…]
LLM2CLIP 15M 91.3 80.6 67.0 50.6 96.3 95.3 86.4 90.5 86.4 88.5 71.7 75.3 83.2 80.1
[…]
LLM2CLIP 30M 93.1 81.0 68.2 52.0 97.5 97.7 92.7 93.9 88.2 89.6 74.9 78.3 85.8 82.1
[…]
Table 1: Cross-modal retrieval performance Recall@1 on multiple datasets.
Figure 3: Per-language image-text retrieval performance on the XM3600 benchmark.
where $\tau$ is a learnable temperature parameter. Beyond standard contrastive learning, we impose a self-distillation constraint on the CLIP image encoder to mitigate excessive forgetting of pre-trained knowledge during adaptation—essential for preserving generalization. On the image encoder side, we apply a regularization loss that is symmetric to the one used in the first stage(Eq. 2, Eq. 3):
$\mathcal{L}_{reg}=\sum_{i=1}^{\mathcal{B}}\mathcal{L}_{ins}(v_{i},v_{i}^{*})+\sum_{i<j}^{\mathcal{B}}\mathcal{L}_{struct}(v_{i},v_{j}^{*})$ (5)
where $v_{i}^{*}=\mathcal{E}_{I}^{*}(I_{i}),v_{j}^{*}=\mathcal{E}_{I}^{*}(I_{j})$ denote the vision embeddings derived from the EMA-updated image encoder. The EMA update procedure is defined as follows:
$\mathcal{E}_{I}^{*}=\alpha\mathcal{E}_{I}^{*}+(1-\alpha)\mathcal{E}_{I},$ (6)
where $\alpha$ controls the update rate of the teacher model parameters. The overall loss function of the contrastive tuning stage is defined as $\mathcal{L}_{\text{tune}}=\mathcal{L}_{\text{info}}+\lambda_{reg}\mathcal{L}_{\text{reg}}$, where $\lambda_{reg}$ is a loss weight.
4 Experiments
Method #Data
Food101
CIFAR10
CIFAR100
SUN397
Cars
Aircraft
DTD
Pets
Caltech101
Flowers
ImageNet
Avg.
Model Architecture: CLIP ViT-B/32
CLIP 400M 83.1 88.7 63.5 61.5 57.6 18.8 42.8 84.6 89.4 66.0 61.9 65.2
LLM2CLIP 3M 49.6 89.2 61.5 60.3 11.5 8.6 47.8 38.0 79.0 22.6 41.0 46.3
[…]
LLM2CLIP 15M 57.2 88.3 61.4 61.3 19.6 8.4 50.6 42.3 80.7 23.5 43.3 48.8
[…]
LLM2CLIP 30M 58.5 88.3 61.0 61.2 20.6 8.4 50.3 37.6 81.7 26.0 45.1 49.0
[…]
Model Architecture: CLIP ViT-B/16
CLIP 400M 87.9 89.7 66.8 63.1 63.7 22.8 45.0 87.0 90.4 67.6 67.1 68.3
LLM2CLIP 3M 56.9 92.6 64.4 62.2 15.4 11.7 50.9 46.5 82.9 23.6 45.8 50.3
[…]
LLM2CLIP 15M 63.2 90.8 64.5 62.9 27.3 9.9 52.8 50.3 83.2 23.7 46.5 52.3
[…]
LLM2CLIP 30M 64.4 90.2 64.6 63.7 27.0 11.2 55.0 45.9 84.0 27.1 49.7 53.0
[…]
Model Architecture: CLIP ViT-L/14
CLIP 400M 92.6 94.9 77.0 66.8 76.5 30.7 54.4 93.2 93.9 78.1 74.5 75.7
LLM2CLIP 3M 64.8 95.4 72.9 66.4 18.8 10.4 54.8 47.3 88.3 26.8 52.8 54.4
[…]
LLM2CLIP 15M 70.1 95.2 72.3 66.4 32.4 9.5 58.0 54.3 88.3 26.6 54.0 57.0
[…]
LLM2CLIP 30M 71.2 94.0 70.5 67.0 32.1 11.3 57.8 54.7 89.3 28.8 56.4 57.5
[…]
Table 2: Zero-shot classification performance on 11 datasets. The best results are marked in bold.
Method #Data Robustness
IN-V2 IN-A IN-O IN-R IN-S
Model Architecture: CLIP ViT-L/14
CLIP 400M 69.8 70.8 32.2 87.8 59.6
LLM2CLIP 3M 49.0 46.6 32.4 75.0 44.8
[…]
LLM2CLIP 15M 50.8 50.1 33.8 78.2 46.3
[…]
LLM2CLIP 30M 52.7 52.7 34.0 78.6 47.3
[…]
Table 3: Robustness performance. The best results are marked in bold.
4.1 Experimental Setup
Datasets and Benchmarks. For the alignment dataset, we use CC3M (Changpinyo et al., 2021), CC12M (Changpinyo et al., 2021), and YFCC15M (Thomee et al., 2016),combined the high-quality captions from DreamLIP (Zheng et al., 2024). We conduct experiments with data scales of 3M (CC3M), 15M (CC3M + CC12M), and 30M (CC3M + CC12M + YFCC15M) to explore the effects of data scaling. For the benchmark, we perform zero-shot classification on 11 different classification datasets, robustness evaluations on 5 ImageNet variants, retrieval evaluations on 6 datasets, multilingual cross-modal retrieval evaluation on XM3600 (Thapliyal et al., 2022), and fine-grained understanding evaluation on MMVP-VLM (Tong et al., 2024) and SugarCrepe (Hsieh et al., 2023). Regarding the model, we mainly employ three OpenAI pre-trained CLIP models, ViT-B/32, ViT-B/16, and ViT-L/14, to investigate the effects of model scaling. Additionally, we conduct experiments with pretrained EVA02-CLIP (Fang et al., 2023) ViT-L/14 and SigLIP2 Tschannen et al. (2025) SO/14@224 to assess the impacts of different model architectures. For the LLM-based embedder, we primarily use LLaMA3-8B-CC consistent with LLM2CLIP Huang et al. (2024).
Figure 4: Compositional performance on SugarCrepe.
Implementation Details. For the representation inheritance phase, we train for four epochs, followed by another four epochs for contrastive tuning. During training, we employ AdamW (Loshchilov, 2019) as the optimizer, with a learning rate of $1\times 10^{-5}$ and a weight decay of 0.2. The parameters $\beta_{1}$ and $\beta_{2}$ are set to 0.9 and 0.98, respectively. In the first stage, the training batch size is set to 1024, while in the second stage it is increased to 4096. The loss weight $\lambda_{r}eg$ is set at 0.0004. Other training details can be found in the appendix.
Figure 5: MMVP performance.
Method ViT Init LLM Embedder #Data IN-1k COCO Flickr30k
I2T T2I I2T T2I
FLAME random Mistral-Nemo 3M 36.0 43.3 28.6 67.3 53.6
ShareLock DINOv2 B/14 Llama3 3M 52.1 - - - -
LIFT random NV-Embedv2 512M 43.6 34.6 36.0 69.1 72.9
LiT CLIP B/16 Llama3-$C$ 3M 51.0 56.2 41.9 85.2 71.9
LLM2CLIP CLIP B/16 Llama3-$C$ 3M 45.8 60.5 44.8 88.0 75.3
[…]
SAIL DINOv2 L/14 NV-Embedv2 3M 54.0 45.4 32.9 - -
LiT CLIP L/14 Llama3-$C$ 3M 60.1 59.4 44.6 88.0 74.7
LLM2CLIP CLIP L/14 Llama3-$C$ 3M 52.8 65.5 49.7 92.4 80.1
[…]
Figure 6: Comparison with other methods across different model scales and LLM embedders.
4.2 Main Results
Cross-Modal Retrieval. Tab. 1 shows that ProCLIP consistently outperforms LLM2CLIP across all datasets and model scales. In short-text scenarios (e.g., Flickr30k and COCO), ProCLIP achieves substantial gains; notably, with ViT-L/14 (30M), it reaches 95.0% I2T R@1 on Flickr30k, outperforming LLM2CLIP by  2%. This advantage extends to long-text benchmarks (DOCCI, DCI, Urban-1k), where ProCLIP (ViT-B/16, 30M) attains 73.6% I2T and 76.9% T2I R@1 on DCI. Across all training scales (3M to 30M), ProCLIP yields stable improvements, particularly in T2I retrieval. These results validate ProCLIP’s versatility in handling both concise and complex textual scenarios.
Multilingual cross-modal Retrieval. Leveraging the LLM-based embedder, ProCLIP exhibits robust multilingual capabilities. As illustrated in Fig. 3, ProCLIP consistently outperforms LLM2CLIP on the XM3600 benchmark (Thapliyal et al., 2022) across various languages. This superiority is primarily attributed to our progressive alignment strategy, which more effectively bridges the semantic gap between the CLIP image encoder and the multilingual LLM space while preserving pretrained visual priors.
Zero-Shot Classification. Tab. 2 summarizes zero-shot classification performance across 11 downstream tasks. We observe that LLM2CLIP severely compromises CLIP’s inherent generalization, with average performance dropping by 16.2% (ViT-B/32), 15.3% (ViT-B/16), and 18.2% (ViT-L/14) even when trained on 30M samples. In contrast, ProCLIPconsistently achieves substantial improvements over LLM2CLIP across all configurations, yielding a 10.0%–13.5% average gain on the 30M dataset. This superiority stems from the well-crafted alignment curriculum of ProCLIP, which leverages pretrained knowledge to achieve more effective and robust cross-modal alignment.
Robustness. Tab. 3 evaluates the robustness of ProCLIP across various data scales and model architectures, showing consistent average gains of 5.9%–9.3%. Notably, on challenging out-of-distribution (OOD) benchmarks such as ImageNet-A and ImageNet-R, ProCLIP outperforms LLM2CLIP by over 10 percentage points. This substantial margin highlights ProCLIP’s superior capacity to handle distribution shifts and complex perturbations.
Figure 7: Ablation on different LLM-based embedders.
Method Stage 1 Stage 2 IN-1k I2T Avg. T2I Avg.
$\mathcal{L}_{\text{ins}}$ $\mathcal{L}_{\text{struct}}$ $\mathcal{L}_{\text{info}}$ $\mathcal{L}_{\text{reg}}$
CLIP 74.5 66.9 55.5
LLM2CLIP 52.8 82.0 79.1
ProCLIP 58.9 69.3 59.4
59.5 70.3 61.2
59.2 82.9 80.2
[…]
Figure 8: Ablation on different components.
Fine-Grained Understanding. Fig. 4 demonstrates the significant potential of ProCLIP in compositional understanding tasks, surpassing the CLIP baseline by an average margin of 11.3%–17.7%. These improvements demonstrate that the LLM-based embedder enhances fine-grained semantic discrimination, and the consistent superiority of our method underscores the effectiveness of the progressive alignment strategy. Fig. 6 presents the fine-grained vision-language understanding performance on the MMVP benchmark (Tong et al., 2024) using CLIP ViT-L/14. LLM2CLIP improves over CLIP by 2.9%, 5.9%, and 4.4% at 3M, 15M, and 30M data scales, respectively. Our ProCLIP model further advances these results, achieving gains of 3.0%, 2.2%, and 10.4% on the corresponding data scales.
Comparison with Other Methods. To further prove the effectiveness of ProCLIP, we provide a comprehensive comparison of all recent LLM embedder-based CLIP models, including FLAME, ShareLock, LIFT, SAIL, LiT, and our baseline LLM2CLIP. As shown in Tab. 6, under the same or lower training costs, ProCLIP consistently achieves superior performance across various model sizes. Benefiting from representation inheritance and self-distillation regularization, ProCLIP not only achieves significant performance improvements in In-1k classification but also enhances general retrieval capabilities on COCO and Flickr30k.
4.3 Ablation Study
Ablation of Different LLM-based Embedder. Fig. 8 compares several LLM embedders (Qwen3, GME, NV-Embedv2, and Llama3-CC) paired with ViT-L/14. Among them, Llama3-CC achieves the best overall results in classification and retrieval. Notably, while retrieval performance remains relatively consistent across embedders, ImageNet classification accuracy varies significantly. This suggests that the alignment discrepancy between each LLM’s feature space and the original CLIP space differs, leading to varying degrees of generalization degradation after training.
Base Model Method IN-1k I2T Avg. T2I Avg.
CLIP ViT-L LLM2CLIP 52.8 82.0 79.1
[…]
EVA02-CLIP ViT-L LLM2CLIP 56.4 83.6 80.9
[…]
SigLIP2 SO/14@224 LLM2CLIP 63.8 84.6 81.9
[…]
Table 4: Ablation study on different base model architectures. All methods use 3M training samples.
Ablation of Different Base Model. Tab. 4 shows ProCLIP excels across architectures, surpassing LLM2CLIP in classification and retrieval. The remaining classification gap relative to vanilla CLIP arises from the latter’s overwhelming data scale, a dominant factor in zero-shot tasks.
Ablation of $\lambda_{reg}$. Tab. 5 presents the sensitivity analysis of the regularization weight for $\lambda_{\text{reg}}$. The results indicate that 4e-4 offers an optimal balance and is thus selected as our final configuration. Furthermore, the empirical evidence suggests that excessively strong regularization hampers alignment, leading to suboptimal results.
Ablation of Different Components. Tab. 8 details the ablation study of our components. First, instance semantic distillation achieves 58.9% ImageNet-1k zero-shot accuracy using text only, confirming the successful transfer of CLIP’s textual priors to the adapter. Adding structural alignment loss further boosts both classification and retrieval by enabling the LLM to capture CLIP’s global geometric structure beyond point-wise semantics. Subsequent image-text contrastive tuning significantly enhances retrieval but compromises classification due to image encoder overfitting. This is mitigated by self-distillation, which recovers classification accuracy from 59.2% to 62.5%. Ultimately, structured self-distillation yields the best results by stabilizing the image representation space and preserving pretrained knowledge, effectively balancing fine-tuning gains with generalization.
$\lambda_{reg}$ IN-1k I2T Avg. T2I Avg.
0 59.2 82.9 80.2
2e-4 60.3 82.8 80.3
[…]
6e-4 62.6 82.6 79.8
1e-3 62.5 81.8 79.5
Table 5: Ablation of $\mathcal{L}_{reg}$
5 Conclusion
This paper presents ProCLIP, a progressive vision-language alignment framework for integrating CLIP image encoders with LLM-based embedders. Inspired by curriculum learning, ProCLIP employs a two-stage strategy: first, it anchors the LLM-based embedder to CLIP’s textual space via knowledge distillation to inherit pretrained semantic priors; second, it performs cross-modal contrastive tuning while utilizing self-distillation to prevent overfitting. To maintain feature-space consistency, we introduce a complementary distillation scheme incorporating instance semantic and structural alignment losses. Extensive evaluations across various scales and architectures validate the efficacy and generality of our approach.
6 Limitation
[…]
Appendix A Appendix
A.1 Training Details
Details of the hyperparameter configurations used for two-stage training of ProCLIP are presented in Tab. 6. Under the default setting, our MLP layers are consistent with the baseline LLM2CLIP, both consisting of four linear layers.
Hyperparameters of stage1
Batch size 1024 ($8\times 128$)
Optimizer AdamW
Weight decay 0.05
Adam $\beta$ (0.9,0.98)
Adam $\epsilon$ 1e-6
Learning rate 1e-5
Learning rate schedule cosine decay
Epochs 4
Training GPUs $8\timesH$100
Hyperparameters of stage2
Batch size 4096 ($8\times 512$)
Optimizer AdamW
Weight decay 0.05
Adam $\beta$ (0.9,0.98)
Adam $\epsilon$ 1e-6
Learning rate 1e-5
Learning rate schedule cosine decay
Ema $\alpha$ 0.999
$\lambda_{\text{reg}}$ 0.0004
Epochs 4
Training GPUs $8\timesH$100
Table 6: Detailed hyperparameters for training ProCLIP.
A.2 Details of Benchmarks.
Zero-Shot Classification & Linear Probe. Following the previous works (Yang et al., 2023; Gu et al., 2024), we evaluate the zero-shot classification and linear probe performance of the models on 11 datasets. The detailed information about these datasets and the prompt used in zero-shot classification are presented in Tab. 7 and Tab. 14.
Dataset Classes Train size Test size Evaluation metric
Food101 102 75,750 25,250 accuracy
CIFAR10 10 50,000 10,000 accuracy
CIFAR100 100 50,000 10,000 accuracy
SUN397 397 19,850 19,850 accuracy
Cars 196 8,144 8,041 accuracy
Aircraft 100 6,667 3,333 mean per class
DTD 47 3,760 1,880 accuracy
Pets 37 3,680 3,669 mean per class
Caltech101 101 3,000 5,677 mean-per-class
Flowers 102 2,040 6,149 mean per class
ImageNet 1000 1,281,167 50,000 accuracy
Table 7: List of zero-shot datasets with the data distribution and evaluation metrics.
Robustness. We evaluated the robustness of our model on five out-of-distribution datasets, including ImageNet-v2 (Recht et al., 2019), ImageNet-A (Hendrycks et al., 2021b), ImageNet-O (Hendrycks et al., 2021b), ImageNet-R (Hendrycks et al., 2021a), and ImageNet-Sketch (Wang et al., 2019).
Cross-Modal Retrieval. Following the previous works (Huang et al., 2024; Cao et al., 2025), we evaluate the cross-modal retrieval performance of the models on 6 datasets: Flickr30k (Plummer et al., 2015), COCO (Lin et al., 2014), ShareGPT4V (Chen et al., 2024), Urban-1k (Zhang et al., 2024), DOCCI (Onoe et al., 2024), and DCI (Urbanek et al., 2024). The details information about these dataset are present on Tab. 8.
Dataset Test Images Evaluation Protocol Text type
MSCOCO 5,000 Image-to-Text & Text-to-Image short
Flickr30k 1,000 Image-to-Text & Text-to-Image short
ShareGPT4V 1000 Image-to-Text & Text-to-Image long
Urban-1k 1000 Image-to-Text & Text-to-Image long
DOCCI 5000 Image-to-Text & Text-to-Image long
DCI 7805 Image-to-Text & Text-to-Image long
Table 8: Zero-shot image-text retrieval evaluation settings.
Multilingual Retrieval. We evaluated the multilingual capabilities of our model on XM3600 (Thapliyal et al., 2022). XM3600 contains 3,600 images covering a total of 36 languages, including Arabic (ar), Bengali(bn), Chinese-Simplified (zh), Croatian (hr), Czech (cs), Danish (da), Dutch (nl), English (en),Farsi (fa), Filipino (fil), Finnish (fi), French (fr), German (de), Greek (el), Hebrew (he), Hindi (hi), Hungarian (hu), Indonesian (id), Italian (it), Japanese (ja), Korean (ko),Maori(mi), Norwegian (no), Persian (fa), Polish (pl), Portuguese (pt), Romanian (ro), Russian (ru), Spanish (es), Swedish (sv), Swahili(sw), Thai (th), Turkish (tr), Telugu (te), Ukrainian (uk), and Vietnamese (vi).
Fine-Grained Understanding. We evaluated the fine-grained understanding capability of the VLM on MMVP-VLM (Tong et al., 2024) and SugarCrepe. MMVP-VLM consists of 150 samples in total, testing 9 patterns:
REPLACE SWAP ADD
Object Attribute Relation Object Attribute Object Attribute
# negative captions 1652 788 1406 246 666 2062 692
Table 9: Number of hard negative captions of all types in SugarCrepe.
Method Flickr30k COCO ShareGPT4V Urban-1k DOCCI DCI Avg.
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
CLIP 86.6 64.6 57.2 36.4 78.0 68.7 68.4 56.0 65.8 63.1 45.4 43.9 66.9 55.5
Long-CLIP 89.7 75.7 62.5 46.4 93.3 92.1 82.4 86.2 66.4 78.6 57.2 64.0 75.2 73.8
[…]
Table 10: Compared to Long-CLIP on multiple cross-modal retrieval datasets.
  • Orientation and Direction: Questions about the direction something is facing or moving, such as the direction the dog or duck is facing, or the orientation of the school bus.
  • Presence of Specific Features: Questions that focus on the existence or non-existence of certain elements or features in the image.
  • State and Condition: Questions that pertain to the state or condition of an object, such as whether a flag is blowing in the wind or if the ground is wet.
  • Quantity and Count: Questions about the number of objects or features present in the image.
  • Positional and Relational Context: This aspect refers to the model’s ability to understand the position and relationship of objects or elements within an image in relation to each other and their surroundings.
  • Color and Appearance: Questions regarding the color of certain objects or elements.
  • Structural and Physical Characteristics: This category involves the model’s ability to identify and analyze the physical attributes and structural features of objects in an image.
  • Text: Questions related to text or symbols present in the image.
  • Viewpoint and Perspective: Questions concerning the perspective from which the photo was taken.
SugarCrepe contains three types of negative samples—Add, Replace, and Swap—designed to probe whether vision-language models (VLMs) can recognize text that differs only in subtle structural changes yet conveys a completely different overall meaning in image-text matching tasks.
MLLM benchmarks. We further integrate the fine-tuned vision encoder into LLaVA and evaluate its performance on several MLLM downstream benchmarks, including SEED-Bench (Li et al., 2023a), GQA (Hudson and Manning, 2019), VizWiz (Gurari et al., 2018), PoPE (Li et al., 2023b), TextVQA (Singh et al., 2019), MMBench (Liu et al., 2024b), and VQAv2 (Goyal et al., 2017).
Method SEED-Bench (image) GQA VizWiz PoPE TextVQA MMBench VQAv2
CLIP 65.3 62.0 44.0 85.7 54.2 65.5 77.4
LLM2CLIP 66.4 61.7 44.6 86.3 55.0 65.5 77.9
[…]
Table 11: MLLM(7B) performance under $224^{2}$ image resolution.
Method Data Adapter IN-1k I2T Avg. T2I Avg.
CLIP 400M - 74.5 66.9 55.5
[…]
LLM2CLIP 3M $4\timeslinear$ 52.8 82.0 79.1
[…]
LLM2CLIP 15M $4\timeslayers$ 54.0 83.2 80.1
[…]
LLM2CLIP 30M $4\timeslinear$ 56.4 85.8 82.1
[…]
Table 12: Comparison of data and model scales under ViT-L architecture.
A.3 More results.
Compared to Long-CLIP. Long-CLIP Zhang et al. (2024) is a classic method for extending the long-text capabilities of CLIP. Here, we compare ProCLIP with Long-CLIP. We restricted the data volume for ProCLIP to 1M, consistent with the Long-CLIP experiments. The results are shown in Table 10. Our method outperforms the Long-CLIP paradigm in both short-text and long-text retrieval. Furthermore, we extend CLIP’s capabilities to be both multilingual and fine-grained understanding.
Linear Probe. We conduct linear probe evaluations of the model on 11 datasets. As shown in Tab.7, our method consistently achieves superior performance. This advantage stems from our progressive alignment framework, which stabilizes training through two-stage regularization that prevents overfitting in the vision encoder while preserving generalization capability.
Further Analysis of Data Scale and Model Scale. We further analysis the effects of data scale and model scale, as shown in Tab. 12. For data scale, we observe that model performance improves with increasing data size. For example, when trained on 3M samples, ProCLIP achieves a zero-shot IN-1k accuracy of 62.5, which rises to 67.8 when the dataset size increases to 30M. Under the same data scale, ProCLIP consistently outperforms LLM2CLIP. Notably, when we randomly sample 1M images from CC3M for training, ProCLIP achieves comparable or even superior zero-shot retrieval performance compared with LLM2CLIP, reaching 61.8 on zero-shot IN-1k. This highlights the data efficiency of ProCLIP. For model scale, we further expand the linear layers by three times, using 12 layers in total, which leads to additional performance gains. This suggests that ProCLIP can continue to benefit from simple parameter scaling.
MLLM Performance. As shown in Tab. 11, when integrating the fine-tuned vision encoder into the MLLM, we observe performance improvements over CLIP on most benchmarks. This can be mainly attributed to the alignment with high-quality data, which enhances the semantic representation capability of the vision encoder. ProCLIP and LLM2CLIP achieve relatively comparable performance, indicating that ProCLIP does not exhibit a significant advantage within the MLLM benchmarks. We attribute this to the fact that our method, compared with the baseline, does not place additional emphasis on the downstream MLLM benchmarks. A further discussion on this issue can be found in Openvision Li et al. (2025) and OpenVision2 (Liu et al., 2025).
Method Data
Food101
CIFAR10
CIFAR100
SUN397
Cars
Aircraft
DTD
Pets
Caltech101
Flowers
ImageNet
Avg.
Model Architecture: CLIP ViT-B/32
CLIP 400M 88.6 95.1 80.1 73.4 80.8 44.9 76.3 89.3 92.7 94.7 74.3 80.9
LLM2CLIP 3M 87.9 95.7 83.1 74.1 78.0 44.9 77.7 90.4 92.4 94.6 74.2 81.2
[…]
LLM2CLIP 15M 87.7 95.7 82.7 74.0 77.5 44.2 78.3 90.2 92.5 94.4 74.2 81.0
[…]
LLM2CLIP 30M 87.6 95.9 83.0 74.1 76.3 43.5 77.6 90.1 92.8 93.8 74.3 80.8
[…]
Model Architecture: CLIP ViT-B/16
CLIP 400M 92.7 96.0 82.5 75.7 85.9 52.8 78.9 93.1 93.9 96.4 79.6 84.4
LLM2CLIP 3M 91.6 97.0 84.5 76.0 82.1 50.1 80.3 92.3 93.6 95.7 79.6 83.9
[…]
LLM2CLIP 15M 91.9 97.0 84.9 75.6 83.7 50.7 80.4 92.9 93.8 96.6 79.6 84.3
[…]
LLM2CLIP 30M 91.3 96.6 84.8 75.3 80.6 48.2 80.3 92.5 93.4 95.0 79.7 83.4
[…]
Model Architecture: CLIP ViT-L/14
CLIP 400M 95.3 89.1 87.2 79.4 90.7 63.0 81.8 95.3 96.9 98.8 82.9 88.1
LLM2CLIP 3M 94.5 98.6 89.2 79.6 86.7 57.7 83.4 94.1 96.4 97.1 82.5 87.2
[…]
LLM2CLIP 15M 94.4 98.5 88.8 78.5 86.0 55.0 82.7 93.9 95.9 97.1 82.6 86.7
[…]
LLM2CLIP 30M 94.1 98.2 88.4 78.7 84.6 54.8 82.4 93.7 95.8 96.5 82.2 86.3
[…]
Model Architecture: EVA02-CLIP ViT-L/14
EVA02-CLIP 2B 95.6 99.5 94.2 80.4 94.2 69.5 85.0 94.8 97.6 99.4 84.1 90.4
LLM2CLIP 3M 94.1 99.5 93.3 79.4 85.0 54.3 84.0 93.2 97.3 96.9 84.1 87.4
[…]
Table 13: Linear Probe performance on 11 datasets.
CIFAR 10 & CIFAR 100
a photo of a {label}. a blurry photo of a {label}. a black and white photo of a {label}. a low contrast photo of a {label}.
a high contrast photo of a {label}. a bad photo of a {label}. a good photo of a {label}. a photo of a small {label}.
a photo of a big {label}. a photo of the {label}. a blurry photo of the {label}. a black and white photo of the {label}.
a low contrast photo of the {label}. a high contrast photo of the {label}. a bad photo of the {label}. a good photo of the {label}.
a photo of the small {label}. a photo of the big {label}.
Food101
a photo of {label}, a type of food.
Caltech101
a photo of a {label}. a painting of a {label}. a plastic {label}. a sculpture of a {label}.
a sketch of a {label}. a tattoo of a {label}. a toy {label}. a rendition of a {label}.
a embroidered {label}. a cartoon {label}. a {label} in a video game. a plushie {label}.
an origami {label}. art of a {label}. graffiti of a {label}. a drawing of a {label}.
a doodle of a {label}. a photo of the {label}. a painting of the {label}. the plastic {label}.
a sculpture of the {label}. a sketch of the {label}. a tattoo of the {label}. the toy {label}.
a rendition of the {label}. the embroidered {label}. the cartoon {label}. the {label} in a video game.
the plushie {label}. the origami {label}. art of the {label}. graffiti of the {label}.
a drawing of the {label}. a doodle of the {label}.
Stanford Cars
a photo of a {label}. a photo of the {label}. a photo of my {label}. i love my {label}!
a photo of my dirty {label}. a photo of my clean {label}. a photo of my new {label}. a photo of my old {label}.
DTD
a photo of a {label} texture. a photo of a {label} pattern. a photo of a {label} thing. a photo of a {label} object.
a photo of the {label} texture. a photo of the {label} pattern. a photo of the {label} thing. a photo of the {label} object.
FGVC Aircraft
a photo of a {label}, a type of aircraft. a photo of the {label}, a type of aircraft.
Flowers102
a photo of a {label}, a type of flower.
Pets
a photo of a {label}, a type of pet.
SUN39
a photo of a {label}. a photo of the {label}.
ImageNet
a bad photo of a {label}. a photo of many {label}. a sculpture of a {label}. a photo of the hard to see {label}.
a low resolution photo of the {label}. a rendering of a {label}. graffiti of a {label}. a bad photo of the {label}.
a cropped photo of the {label}. a tattoo of a {label}. the embroidered {label}. a photo of a hard to see {label}.
a bright photo of a {label}. a photo of a clean {label}. a photo of a dirty {label}. a dark photo of the {label}.
a drawing of a {label}. a photo of my {label}. the plastic {label}. a photo of the cool {label}.
a close-up photo of a {label}. a black and white photo of the {label}. a painting of the {label}. a painting of a {label}.
a pixelated photo of the {label}. a sculpture of the {label}. a bright photo of the {label}. a cropped photo of a {label}.
a plastic {label}. a photo of the dirty {label}. a jpeg corrupted photo of a {label}. a blurry photo of the {label}.
a photo of the {label}. a good photo of the {label}. a rendering of the {label}. a {label} in a video game.
a photo of one {label}. a doodle of a {label}. a close-up photo of the {label}. a photo of a {label}.
the origami {label}. the {label} in a video game. a sketch of a {label}. a doodle of the {label}.
an origami {label}. a low resolution photo of a {label}. the toy {label}. a rendition of the {label}.
a photo of the clean {label}. a photo of a large {label}. a rendition of a {label}. a photo of a nice {label}.
a photo of a weird {label}. a blurry photo of a {label}. a cartoon {label}. art of a {label}.
a sketch of the {label}. a embroidered {label}. a pixelated photo of a {label}. itap of the {label}.
a jpeg corrupted photo of the {label}. a good photo of a {label}. a plushie {label}. a photo of the nice {label}.
a photo of the small {label}. a photo of the weird {label}. the cartoon {label}. art of the {label}.
a drawing of the {label}. a photo of the large {label}. a black and white photo of a {label}. the plushie {label}.
a dark photo of a {label}. itap of a {label}. graffiti of the {label}. a toy {label}.
itap of my {label}. a photo of a cool {label}. a photo of a small {label}. a tattoo of the {label}.
Table 14: Full list of prompts to evaluate the performance of zero-shot classification on 11 visual recognition datasets.
A.4 Future Works.
Training Efficiency. ProCLIP is subject to the limitation of increased computational overhead. We consider the following directions to potentially reduce computational overhead:
  • Adopting a PEFT-based approach to fine-tune the vision encoder in the second stage
  • Fine-tuning only part of the vision encoder parameters in the second stage, such as the last few Transformer blocks
  • Replacing online distillation with offline distillation, which would substantially reduce the additional computational cost introduced in the second stage
Fine-grained Visual Alignment. Future work explores the integration of a dedicated local alignment loss to explicitly strengthen the model’s fine-grained visual-textual alignment capabilities, thereby enhancing performance on tasks that require detailed spatial understanding.
More Model Architecture. Our approach replaces the original CLIP text encoder with an LLM-based embedder to enhance multiple capabilities. From another perspective, we consider whether the vision encoder in the dual-tower architecture can also be replaced to address limitations in visual representation. For instance, the CLIP image encoder is known to lack locality; substituting the image encoder may mitigate this limitation. We leave a thorough investigation of this direction for future work.
A.5 Visualization.
As illustrated in Fig. 9, ProCLIP effectively extends long-text processing capabilities, enabling precise image-text alignment even within highly detailed descriptions.
Figure 9: The visualization of ProCLIP vs CLIP in long-text cross-modal retrieval. From top to bottom: Case 1, Case 2, Case 3, and Case 4.